OpenStudy (anonymous):

Passes through (1,4) and is parallel to y=-4x-5, Now write down the equation of the line.

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

first we need to find the slope of the equation given. The reason we need to do this is because a parallel line will have the same slope. Can you tell me the slope of y = -4x - 5 ? Do you know how to find it ?

OpenStudy (anonymous):

well -4 is the slope i think and -5 is the y intercept if im not mistaken

OpenStudy (texaschic101):

you are correct...the slope is -4 and the y intercept is -5. Now we can use y = mx + b form slope(m) = -4 now use your points..(1,4)...x = 1 and y = 4 now we sub and since we alredy know the slope will be -4, we need to find the y intercept (b). y = mx + b 4 = -4(1) + b 4 = -4 + b 4 + 4 = b 8 = b so your parallel equation is : y = -4x + 8 understand ? Any questions ?

OpenStudy (anonymous):

isn't b suppose to be -5?

OpenStudy (anonymous):

Oh wait, i DO get it throw in the point. Thanks :D

OpenStudy (texaschic101):

that is your original line...a parallel line will contain the same slope but not the same y intercept. The y intercept in the original equation does not matter